# compat=1.1.0

### v1

@@ 8 Acknowledgment
where $H_{\rm inf}$ is the Hubble rate during inflation, $t_{\rm end},t_{\rm rh}$ respectively denote the time at the end of inflation and at the reheating, $a_{0}$ is the scale factor at present, and $H_{0}=67$ km/s/Mpc is the Hubble rate at present [23]. If the entropy is conserved from $t=t_{\rm rh}$ to the present, we have $a_{0}/a(t_{\rm rh})=(g_{\rm eff}T_{R}^{3}/g_{S,{\rm eff},0}T_{0}^{3})^{1/3}$, where $T_{R}$ is the reheating temperature, $g_{\rm eff}$ is the effective relativistic degree of freedom at the reheating, $T_{0}=2.73$ K is the CMB temperature at present, and $g_{S,{\rm eff,now}}=43/11$ is the effective entropy degree of freedom at present. By inserting the above values, the condition for solving the horizon problem is recast into

### v2

@@ 6 Numerical analysis
We discuss the reheating temperature and its implications. For the above benchmark, the reheating temperature $T_{R}$ is found to be
$T_{R}=3.4\times 10^{8}~{}{\rm GeV}.$ (47)
With such a high reheating temperature, the gravitino problem [24, 25] can occur. However, severity of the gravitino problem depends on the soft SUSY breaking mechanism that we do not specify, and so this issue is beyond the scope of the present paper.
Since the reheating temperature is much higher than the soft SUSY breaking masses, the MSSM particles, including the one-particle states of squark fields that have constituted the inflaton $\phi$, reach thermal equilibrium after the reheating. Because the interactions of MSSM particles conserve baryon and lepton numbers, the baryon number yield created from the $\phi$ decay is not washed out by these interactions. The difference between the interactions of MSSM particles that conserve baryon and lepton numbers, and the decay of the inflaton $\phi$ that creates baryon number, comes from the fact that $\phi$ is a superposition of a squark field and an antisquark field. As a result, $\phi$ decays into both a quark$+$gaugino/Higgsino and an antiquark$+$gaugino/Higgsino, and asymmetry in the partial widths of the quark decay mode and the antiquark decay mode induces non-zero baryon number. Conversely, such a superposition state cannot be created from the interactions of MSSM particles, and it should be prepared as an initial condition.
@@ 8 Acknowledgment
where $H_{\rm inf}$ is the Hubble rate during inflation, $t_{\rm end},t_{\rm rh}$ respectively denote the time at the end of inflation and at the reheating, $a_{0}$ is the scale factor at present, and $H_{0}=67$ km/s/Mpc is the Hubble rate at present [26]. If the entropy is conserved from $t=t_{\rm rh}$ to the present, we have $a_{0}/a(t_{\rm rh})=(g_{\rm eff}T_{R}^{3}/g_{S,{\rm eff},0}T_{0}^{3})^{1/3}$, where $T_{R}$ is the reheating temperature, $g_{\rm eff}$ is the effective relativistic degree of freedom at the reheating, $T_{0}=2.73$ K is the CMB temperature at present, and $g_{S,{\rm eff},0}=43/11$ is the effective entropy degree of freedom at present. By inserting the above values, the condition for solving the horizon problem is recast into